OpenStudy (anonymous):

how would I solve 2(X+3)-1=7X ?

OpenStudy (anonymous):

2x+5=7x 5=5x 1=x

OpenStudy (anonymous):

2x+6-1=7x -5x=-5 then x=1

OpenStudy (anonymous):

2x+5=7x then -5x=-5 then x=1

OpenStudy (anonymous):

Jessii, do you understand why the answer is as stated by the others?

OpenStudy (anonymous):

You want to apply ALL the rules you have available to get "X" all by itself on ONE side of the equation.

OpenStudy (anonymous):

oh okay, thank you (:

OpenStudy (anonymous):

It's surprising that I never really learned this - I sort of stumbled on it after many many math classes. Apply the operations or "tools" you know and try to see if you can solve a given problem. The laws / rules /tools are associated law, communicative law, uh, and another one. :-)